Institut for Matematik og Datalogi Odense Universitet

April 2, 1996 JFB

DM11 – F96 – Lecture 9

Lecture, March 27

We covered sections 4.6, 4.7, 5.4 and 5.5.

Lecture, April 10

We will cover sections 5.1 and 5.2.

Lecture, April 17

We will cover section 5.3 and Appendix 3.

Problems to be discussed in week 16

• 4, 6, 8, 16, 18, 19, 22, and 23 from section 5.5 . Note that there is a typo in the textbook in the answer to 23. You may have seen the answer to 23 written in another form. The answer you will

get using this method is equal to the other value: ∏_{i=1}^{m} p_{i}^{a_{i}−1}(p_{i} − 1).

• 11, 16, 17, 31, and 38 from section 5.1.

For number 17, the recurrence relation I found had a different form than the one in the textbook, but it was equivalent. You may also sometimes find different, but equivalent recurrence relations to those given in class or the textbook.

• 3a, 3b, 3c, 3e, 3g, 7, and 16, from section 5.2.

Note: You may turn in number 18 from section 5.5 and 38 from section 5.1 to your instructor for comments.